### v2

@@ Abstract
Including local populations in energy systems planning can increase the legitimacy and socio-political feasibility of energy transitions. Participatory modelling enables comprehensive involvement of stakeholders and can facilitate learning; however, the choice of options and scenarios in the model can be opaque to participants. Here we present a framework in which stakeholders engage with near-optimal modelling results in a holistic way, and showcase our methodology in a case study conducted in the remote Arctic settlement of Longyearbyen. Participants choose across a continuum of modelling-based energy system designs in an interactive interface through which they select essentially any feasible combination of system components, and immediately see the implications of their choice. We observe that stakeholders in Longyearbyen consistently deviate from the cost optimum and instead balance different aspects like emissions, costs, and system vulnerability. When confronted with difficult trade-offs through our tool, participants feel better informed and successfully navigate through complex and intertwined decision-making.
@@
Energy systems modelling, Participation, Near-optimal, Modelling-to-generate-alternatives, Stakeholder involvement
@@ 1 Introduction
Our transitions towards more sustainable energy systems need to accelerate, but face substantial challenges. Such transformations require considerable technological change, development, and innovation, but are also increasingly recognised to rely on changes in human practices [1] as well as societal and citizen engagement [2, 3]. Despite barriers of social acceptance and political feasibility [4], much research analysing future energy systems devotes surprisingly little focus to it [5, 6, 7, 8]. Energy transitions involving people in decision-making create legitimacy of the processes, enables change and can help ensure socially just outcomes [9].
@@
Participatory research in energy modelling aims to fill this gap, but currently lacks holistic and simultaneously engaging methodologies. In this article, we introduce a multi-step framework through which policymakers, stakeholders, and modellers can learn from each other and interact to find more socially acceptable energy transitions; we demonstrate this framework with a case study conducted in the Arctic community of Longyearbyen. Our methodology is rooted in a vast number of near-optimal realistic system designs presented in an interactive interface, allowing study participants to consider trade-offs and communicate priorities for system planning. We introduce a modelling approach, to our knowledge the first of its kind, in which participants can interact with and decide from a vast space of feasible energy systems to design a system of their liking.
@@
Despite generally favourable public perceptions of renewable electricity technologies [10, 11] and transitions to low-carbon energy systems altogether in Europe, opinions vary on the favoured end-state and pathways to get there [12]. The overhaul of energy infrastructure to low-carbon systems comes with environmental [13, 14] as well as social consequences [15, 16, 17]. Although it is crucial to consider such (in-)direct impacts for public acceptance, e.g. querying laypeople and experts, the knowledge gaps about technologies in question can lead to misconceptions of risks and impacts [18, 19] when surveys are used to evaluate different energy system components. Other informative approaches such as factsheets [20], group discussions [18], and interactive tools [20, 21] generate more stable preferences [19], create more knowledge and are additionally appreciated by participants [20, 22, 18]. Yet, in all of the above it is challenging to represent the underlying system holistically, capturing the complex interplay of variable renewables, transmission and storage [23]. This motivates the usage of energy system models whose techno-economic nature gives insights about different transition pathways and which challenges exist. Increasingly, energy modellers engage in participatory modelling: they involve stakeholders already at the modelling stage through scenario generation, surveys and interviews, or multi-criteria decision-making [24]. Particular stakeholder groups may still be overlooked, as only one in six reviewed participatory modelling studies included non-academic stakeholders [24]. Nonetheless, participatory modelling can generate bottom-up knowledge of assumptions and narratives [25, 24], help assess which designs would be socially accepted [26] and facilitate mutual learning for both researchers and involved actors [24]. With increasingly complex models, the success of such approaches hinges on transparency of assumptions and modelling tools [24], but presenting stakeholders with results based on a few pre-selected scenarios and criteria may constrain the diversity of options available, potentially based on normative preconceptions and assumptions by researchers. Instead, modelling frameworks in which stakeholders themselves are allowed to define metrics and explore alternative system configurations are preferred [25, 27].
@@
A systematic exploration of alternative system designs is possible by describing the near-optimal feasible space of a model, also known as modelling-to-generate alternatives (MGA) [28]. Looking beyond only cost-optimal modelling, MGA finds feasible solutions that are only slightly more expensive, but score better on other objectives: it is possible to directly search for more diverse alternatives [29, 30], classify trade-offs [31], find designs that are robust to weather uncertainty [32], or perform better on non-modelled objectives such as inequity [33], political feasibility, energy security concerns or distributional outcomes [34, 35]. Despite the myriad applications of MGA, and the potential for communicating trade-offs and informing policymakers, it has not been applied to involve stakeholders in participatory modelling efforts. Here we bridge this gap by creating a framework through which near-optimal spaces describe the diversity of possible solutions which are accessed and communicated through an interactive interface.
@@
In this work, we investigate how participatory modelling by near-optimal solutions can inform 1) stakeholders about transition possibilities and 2) policymakers about stakeholder’s priorities. We demonstrate the methodology with a holistic view on the move away from coal in the remote Arctic settlement of Longyearbyen, Svalbard. Its remote location and inability to externalise the impact of energy system infrastructure makes it overseeable and particularly suitable for stakeholder engagement. We pre-compute 56,050 feasible energy system configurations with an energy system optimisation model (PyPSA-LYB) leading to diverse outcomes for the community which are accompanied by metrics measuring different qualities of each system design. These configurations are presented to inhabitants of Longyearbyen through a novel interactive interface which allows us to assess the technological preferences, priorities and rationale of the participants while ensuring that system requirements are met at all times. Our implementation, which dynamically updates selectable design parameters in order to restrict the user to only feasible system designs, represents a novel and accessible way to explore and interact with the vast space of options generated by near-optimal methods. Despite being used in a specific geographical context, the conceptual framework and tool can be implemented in other settings.
@@ 2 A new modelling-based interactive planning tool
Figure 1: Overview over Longyearbyen and its current energy system. On the left, the map shows the geographical location of Longyearbyen on the Arctic archipelago of Svalbard as well as its distance from the north of Norway ($\sim$950 km) making an electricity link prohibitively expensive and unrealistic. On the top right, the per capita greenhouse gas emissions for Longyearbyen (and Norway as comparison) are plotted: for the coal-powered CHP power plant (until 2022), and the current diesel-powered CHP power plant [36, 37]. Norwegian emissions are computed based on 2021 data on electricity consumption (https://www.ssb.no/statbank/table/06913/) and its carbon density (https://www.nve.no/energi/energisystem/kraftproduksjon/hvor-kommer-stroemmen-fra/) as well as biomass heating emissions from Arvesen et al. [38]. On the bottom right, the monthly average capacity factors for onshore wind and solar PV are plotted (for the input weather year). Note that these are subject to interannual variability and therefore just one realisation.
@@
Longyearbyen (overview in figure 1) is well-suited for energy systems research for a number of reasons. For one, the energy system of Longyearbyen is an isolated off-grid system, located at 78°N on Svalbard, and a mainland connection ($\sim$950km away) is prohibitively expensive [39]. The external influences on the system are therefore limited, which simplifies modelling, and at the same time makes impacts and trade-offs of different energy technologies directly apparent. Longyearbyen is already transitioning away from its coal-powered history and is of extraordinary geopolitical interest [40] — which keeps energy on the agenda of the local population, and is frequently discussed in local and national media.
@@
Figure 2: Workflow for participatory modelling studies using near-optimal methods. On the left side we suggest an abstract workflow in which policymakers, modellers, and stakeholders collaborate to develop more socially acceptable transitions. For each of these actors, the circles with “P”, “M”, and “S” indicate at which stages they are involved. On the right side, we present a concrete implementation of this workflow for our case study in Longyearbyen, starting with “Energiplan” as in [41].
@@
To engage the local community in energy planning, we developed an interactive interface, which functions as a tool for users to interact with results from an energy system optimisation model, PyPSA-LYB (Methods). The scenarios underlying the model have previously been used for planning the energy transition in Longyearbyen [41] and serve as a baseline for our modelling development. With the transition plan [41] defining the limits and broad scope of modelling assumptions, we explore near-optimal solutions to paint a realistic picture of how alternative energy system designs for Longyearbyen could look like. In particular, we follow Grochowicz et al. [32] to generate the near-optimal feasible space and sample 56,050 different energy system configurations (see Methods).
@@
We engaged with the local population in public spaces between March 11 and 15, 2024; each participant interacting with the interface individually. In the interface, participants explored a continuum of energy system designs, by allocating investment in different system components. Within the feasible bounds (pre-computed in PyPSA-LYB) and a generous cost increase (of up to 125% on top of the optimal costs), participants freely set investment levels for onshore wind power, solar power, imported green fuels, heat storage and hydrogen infrastructure (see Methods). Figure 2 illustrates both the generalised (left) and particular (right) workflow used.
@@
In practice, the level of investments is controlled through sliders, as seen in figure 3a) and restricted by an updated geometry of the near-optimal space relative to the slider location. Apart from the investment levels, the participants were presented with key metrics of their chosen system configuration. These metrics are computed heuristically as needed by the interface, interpolating between pre-computed values for the 56,050 sampled configurations (Methods). This feature, illustrated in figure 3b), adds relevant information and context to the decision process, and keeps track of our research question aimed at (balancing) priorities of local stakeholders. By complementing these quantitative values with possible qualitative preconceptions, the participants saw the impacts of their choices directly and could identify an energy system design best representing their preferences. After submitting their final energy system design, participants were also asked to choose up to three metrics (some chose more than three) that they prioritised for the energy system design — i.e. their stated priorities —, how much additional cost they were willing to accept to improve the representation of their priorities and provide written feedback to the process. Complete details on the interface, as well as the rationale for choosing the particular dimensions and metrics are described in Methods.
@@ 3 Cost-optimality paradigms cannot meet diverse expectations and technological perspectives
The interface outputs showcase the variety of proposed system designs and give insights about the acceptance of technologies, the importance of specific metrics as well as a partial rebuttal of cost-optimal designs, the gold standard of energy system modelling. Through our exemplary work comprising 5% of the population (126 participants, demographics in Methods), we can also infer specific trends and patterns for system planning in Longyearbyen.
Figure 4: Combined strip- and boxplot of submitted results across the five dimensions. Stars represent the cost-optimal investment combination and one million NOK is approximately equivalent to 86,000 EUR or 90,500 USD in 2025. The min-max normalised median absolute deviations (MAD) quantify the distance between the median and the cost-optimum for the five dimensions, and are 0.06, 0.34, 0.11, 0.17 and 0.22, respectively.
None of the technologies have consistent results with interquartile ranges (IQRs) of 10.9–36.8 percent of the decision interval. This is somewhat expected, as the design of the interface as well as the high slack level (125% more expensive systems are “near-optimal”) provide the participants with a high degree of freedom to select their preferred combination of technologies. In addition to the pre-computed 56,050 discrete energy system designs, the participants are also able to pick points anywhere within the feasible space, as the post-processing and interpolation within the interface (Methods) increases the diversity of options.
@@
The resulting energy system designs of the participants show very high slack values (mean 81% and median 91%), raising questions about how well the additional cost is understood by the participants and who would bear that cost. This is striking given that almost half of the participants prioritised electricity prices when designing their system (more below). On the contrary, when the participants were explicitly asked about acceptable additional cost, the results were much lower (mean and median values of 26% and 15% respectively) in the post-study questionnaire. This stark divide between revealed and stated preferences hints at the challenge to make results more tangible, but also shows that participants expect high performance on many priorities including affordability simultaneously.

### v3

@@ Abstract
Including local populations in energy transition planning can increase its legitimacy, socio-political feasibility and, in turn, climate mitigation efficacy. Comprehensive stakeholder involvement in participatory modelling facilitates learning; however, the choice of options and scenarios in the model can be opaque to participants. Here we present a framework in which stakeholders engage with near-optimal modelling results holistically, and showcase our methodology in a case study of the remote Arctic settlement of Longyearbyen, Svalbard. Participants choose across a continuum of modelling-based energy system designs in an interactive interface through which they select essentially any feasible combination of system components, and immediately see the implications of their choice. With Longyearbyen already affected by climate change, stakeholders consistently deviate from the cost optimum and instead balance different aspects like emissions, costs, and system vulnerability. When confronted with difficult trade-offs, participants feel better informed through our tool and successfully navigate through complex and intertwined decision-making.
@@
Energy systems modelling, Energy systems planning, Participation, Near-optimal, Modelling-to-generate-alternatives, Stakeholder involvement
@@ 1 Introduction
Our transitions towards more sustainable energy systems and to mitigate climate change need to accelerate, but face substantial challenges. Such transformations require considerable technological change, development, and innovation, but are also increasingly recognised to rely on changes in human practices [perlaviciutePerspectiveHumanDimensions2021] as well as societal and citizen engagement [chilversSystemicApproachMapping2021, newellNavigatingTensionsRapid2022]. Despite barriers of social acceptance and political feasibility [wustenhagenSocialAcceptanceRenewable2007], much research analysing future energy systems devotes surprisingly little focus to it [sovacoolIntegratingSocialScience2015, trutnevyteSocietalTransformationsModels2019, susserWhyEnergyModels2022, krummModellingSocialAspects2022]. Energy transitions involving people in decision-making create legitimacy of the processes, enables change and can help ensure socially just outcomes [michelsInnovationsDemocraticGovernance2011].
@@
Participatory research in energy modelling aims to fill this gap, but currently lacks holistic and simultaneously engaging methodologies. In this article, we introduce a multi-step framework through which policymakers, stakeholders, and modellers can learn from each other and interact to find more socially acceptable energy transitions. We demonstrate this framework with a case study conducted in Longyearbyen, Svalbard — an Arctic community that already faces severe ramifications of climate change [sydnesLearningCrisis20152021, bradleySvalbardWinterWarming2025, lapointeClimateExtremesSvalbard2024] and is warming twice as fast as the Arctic, and seven times the global average [nordliRevisitingExtendedSvalbard2020]. Our methodology is rooted in a vast number of near-optimal realistic system designs presented in an interactive interface, allowing study participants to consider trade-offs and communicate priorities for system planning. We introduce a modelling approach, to our knowledge the first of its kind, in which participants can interact with and decide from a vast space of feasible energy systems to design a system of their liking, while being informed of their choices’ impacts.
@@
Despite generally favourable public perceptions of renewable electricity technologies [scheerPublicEvaluationElectricity2013, poortingaEuropeanAttitudesClimate2018] and transitions to low-carbon energy systems altogether in Europe, opinions vary on the favoured end-state and pathways to get there [pohjolainenPublicPerceptionsClimate2018]. The overhaul of energy infrastructure to low-carbon systems comes with environmental [giladBiodiversityLineLife2024, giladBiodiversityImpactsNorways2024] as well as social consequences [goforthAirPollutionDisparities2022, sasseLowcarbonElectricitySector2023, weinandExploringTrilemmaCostefficiency2022]. Although it is crucial to consider such (in-)direct impacts for public acceptance, e.g. querying laypeople and experts, the knowledge gaps about technologies in question can lead to misconceptions of risks and impacts [volkenPerspectivesInformedCitizen2018, duboisInformedCitizenPanels2019] when surveys are used to evaluate different energy system components. Other informative approaches such as factsheets [mayerInformedPublicChoices2014], group discussions [volkenPerspectivesInformedCitizen2018], and interactive tools [mayerInformedPublicChoices2014, holzerSwissElectricitySupply2023] generate more stable preferences [duboisInformedCitizenPanels2019], create more knowledge and are additionally appreciated by participants [mayerInformedPublicChoices2014, pidgeonCreatingNationalCitizen2014, volkenPerspectivesInformedCitizen2018]. Yet, in all of the above it is challenging to represent the underlying system holistically, capturing the complex interplay of variable renewables, transmission and storage [grochowiczSpatiotemporalSmoothingDynamics2024]. This motivates the usage of energy system models whose techno-economic nature gives insights about different transition pathways and which challenges exist. Increasingly, energy modellers engage in participatory modelling: they involve stakeholders already at the modelling stage through scenario generation, surveys and interviews, or multi-criteria decision-making [mcgookinParticipatoryMethodsEnergy2021]. Particular stakeholder groups may still be overlooked, as only one in six reviewed participatory modelling studies included non-academic stakeholders [mcgookinParticipatoryMethodsEnergy2021]. Nonetheless, participatory modelling can generate bottom-up knowledge of assumptions and narratives [vageroCanWeOptimise2023, mcgookinParticipatoryMethodsEnergy2021], help assess which designs would be socially accepted [valenzuela-venegasRenewableSociallyAccepted2024] and facilitate mutual learning for both researchers and involved actors [mcgookinParticipatoryMethodsEnergy2021]. With increasingly complex models, the success of such approaches hinges on transparency of assumptions and modelling tools [mcgookinParticipatoryMethodsEnergy2021], but presenting stakeholders with results based on a few pre-selected scenarios and criteria may constrain the diversity of options available, potentially based on normative preconceptions and assumptions by researchers. Instead, modelling frameworks in which stakeholders themselves are allowed to define metrics and explore alternative system configurations are preferred [vageroCanWeOptimise2023, vageroEffectsFairAllocation2024].
@@
A systematic exploration of alternative system designs is possible by describing the near-optimal feasible space of a model, also known as modelling-to-generate alternatives (MGA) [decarolisUsingModelingGenerate2011]. Looking beyond only cost-optimal modelling, MGA finds feasible solutions that are only slightly more expensive, but score better on other objectives: it is possible to directly search for more diverse alternatives [lombardiPolicyDecisionSupport2020, pickeringDiversityOptionsEliminate2022], classify trade-offs [vangreevenbroekTradingRegionalOverall2025], find designs that are robust to weather uncertainty [grochowiczIntersectingNearoptimalSpaces2023], or perform better on non-modelled objectives such as inequity [pedersenUsingModelingAll2023], political feasibility, energy security concerns or distributional outcomes [lauMeasuringExplorationEvaluation2024, lombardiHumanintheloopMGAGenerate2025].
MGA was originally developed to support decision-making of complex and incompletely defined problems [brillModelingGenerateAlternatives1982, brillMGADecisionSupport1990] and has been used for, among other things, land-use planning [brillModelingGenerateAlternatives1982], water resource usage [changUseMathematicalModels1982] and agricultural planning [jeffreyNearlyOptimalLinear1992]. Already when the concept was introduced [brillUseOptimizationModels1979], MGA focused on reaching beyond the limitations of cost-efficiency and social optimality. Citizens and stakeholders may not agree on public goals and how to achieve them, which calls for new methods of using optimisation models within planning processes. Despite the many applications of MGA, and the potential for communicating trade-offs and informing policymakers, it has largely focused on generating alternatives. Rarely have the alternatives been presented to real-world stakeholders and evaluated, even if such an application was presented already in 1990 with an interactive interface and an airline scheduling problem [brillMGADecisionSupport1990] . This is particularly true for energy systems planning, where MGA has only been coupled with synthetic (i.e. assumed) stakeholder preferences based on past interviews [lombardiHumanintheloopMGAGenerate2025].
Previous participatory modelling exercises with interactive interfaces for energy systems planning have used simulation scenarios (my2050 [allenCarbonReductionScenarios2013, demskiEffectsExemplarScenarios2017, pidgeonCreatingNationalCitizen2014]), technology portfolios (Riskmeter [xexakisAreInteractiveWebtools2019, xexakisModelsWrongTrack2020, volkenPerspectivesInformedCitizen2018, holzerSwissElectricitySupply2023, duboisInformedCitizenPanels2019] and the Portfolio-Building Computer Decision Tool [mayerInformedPublicChoices2014]) or spatial generation potential (COLLAGE [flackeInteractivePlanningSupport2017]) as the underlying method for stakeholder interactions. All of these interactive interfaces are based on presenting simplified modelling results, which “cannot aim to fully capture the complexity of electricity capacity planning” (quoting the documentation of the my2050 tool [departmentforenergysecurityandnetzeroCarbonCalculator2024]) such as the system interactions represented by energy system models. Technology portfolios may be informed by energy systems modelling, but results do not include the complementarity of technologies or detailed supply-demand balancing. Similarly, using spatial generation potential only considers technologies individually and annual energy generation potential. Leveraging recent methodological advancements in MGA [lombardiNearoptimalEnergyPlanning2025] facilitates energy systems planning with results that are both feasible and diverse in terms of structure. Unlike the tools outlined above, this means that complex system interactions between variable renewable resources, storage and backup generation are taken into account. Moreover, this means that total system cost (both capital and operational costs) can be accounted for accurately — a feature that is also only possible because the near-optimal solutions are computed using a model with high temporal resolution. Here we bridge these gaps by creating a framework, based on MGA, through which near-optimal spaces describe the diversity of possible solutions which are assessed and communicated through an interactive interface.
@@
In this work, we investigate how participatory modelling by near-optimal solutions can inform 1) stakeholders about transition possibilities and 2) policymakers about stakeholder’s priorities. We demonstrate the methodology with a holistic view on the move away from coal in the remote Arctic settlement of Longyearbyen, Svalbard. Its remote location and inability to externalise the impact of energy system infrastructure makes it overseeable and particularly suitable for stakeholder engagement, and a possible blueprint for other fossil-driven Arctic communities. We pre-compute 56,050 feasible energy system configurations with an energy system optimisation model (PyPSA-LYB) leading to diverse outcomes for the community which are accompanied by metrics measuring different qualities of each system design. These configurations are presented to inhabitants of Longyearbyen through a novel interactive interface which allows us to assess the technological preferences, priorities and rationale of the participants while ensuring that system requirements are met at all times. Our implementation, which dynamically updates selectable design parameters in order to restrict the user to only feasible system designs, represents a novel and accessible way to explore and interact with the vast space of options generated by near-optimal methods. Despite being used in a specific geographical context, the conceptual framework and tool can be implemented in other settings.
@@ 2 A new modelling-based interactive planning tool
Figure 1: Overview over Longyearbyen and its current energy system. On the left, the map shows the geographical location of Longyearbyen on the Arctic archipelago of Svalbard as well as its distance from the north of Norway ($\sim$950 km) making an electricity link prohibitively expensive and unrealistic. On the top right, the per capita greenhouse gas emissions for Longyearbyen (and Norway as comparison) are plotted: for the coal-powered CHP power plant (until 2022), and the current diesel-powered CHP power plant [statisticsnorwaySamfunnsforholdPaSvalbard2023, longyearbyenlokalstyreEnergiomstillingOvergangslosningBP222022]. Norwegian emissions are computed based on 2021 data on electricity consumption (https://www.ssb.no/statbank/table/06913/) and its carbon density (https://www.nve.no/energi/energisystem/kraftproduksjon/hvor-kommer-stroemmen-fra/) as well as biomass heating emissions from Arvesen et al. [arvesenCoolingAerosolsChanges2018]. On the bottom right, the monthly average capacity factors for onshore wind and solar PV are plotted (for the input weather year). Note that these are subject to interannual variability and therefore just one realisation.
@@
Longyearbyen (overview in figure 1) is well-suited for energy systems research for a number of reasons. For one, the energy system of Longyearbyen is an isolated off-grid system, located at 78°N on Svalbard, and a mainland connection ($\sim$950km away) is prohibitively expensive [themaAlternativerFramtidigEnergiforsyning2018]. The external influences on the system are therefore limited, which simplifies modelling, and at the same time makes impacts and trade-offs of different energy technologies directly apparent. Longyearbyen is already transitioning away from its coal-powered history, a centre of climate change research, and of extraordinary geopolitical interest [osthagenMythsSvalbardGeopolitics2024] — which keeps energy on the agenda of the local population, and is frequently discussed in local and national media.
@@
Figure 2: Workflow for participatory modelling studies using near-optimal methods. On the left side we suggest an abstract workflow in which policymakers, modellers, and stakeholders collaborate to develop more socially acceptable transitions. For each of these actors, the circles with “P”, “M”, and “S” indicate at which stages they are involved. On the right side, we present a concrete implementation of this workflow for our case study in Longyearbyen, starting with “Energiplan” as in [grotteEnergiplanLongyearbyenEnergiomstilling2023].
@@
To engage the local community in energy planning, we developed an interactive interface, which functions as a tool for users to interact with results from an energy system optimisation model, PyPSA-LYB (see Methods). The scenarios underlying the model have previously been used for planning the energy transition in Longyearbyen [grotteEnergiplanLongyearbyenEnergiomstilling2023] and serve as a baseline for our modelling development. With the transition plan [grotteEnergiplanLongyearbyenEnergiomstilling2023] defining the limits and broad scope of modelling assumptions, we explore near-optimal solutions to paint a realistic picture of how alternative energy system designs for Longyearbyen could look like. In particular, we follow Grochowicz et al. [grochowiczIntersectingNearoptimalSpaces2023] to generate the near-optimal feasible space and sample 56,050 different energy system configurations (see Methods).
@@
We engaged with the local population in public spaces between March 11 and 15, 2024 (N=126); each participant interacting with the interface individually. In the interface, participants explored a continuum of energy system designs, by allocating investment in different system components. Within the feasible bounds (pre-computed in PyPSA-LYB) and a generous cost increase (an upper limit of 125% on top of the optimal costs), participants freely set investment levels for onshore wind power, solar power, imported green fuels, heat storage and hydrogen infrastructure (see Methods). Figure 2 illustrates both the generalised (left) and particular (right) workflow used.
@@
In practice, the level of investments is controlled through sliders, as seen in figure 3a) and restricted by an updated geometry of the near-optimal space relative to the slider location. Apart from the investment levels, the participants were presented with key metrics of their chosen system configuration. These metrics are computed heuristically as needed by the interface, interpolating between pre-computed values for the 56,050 sampled configurations (see Methods). This feature, illustrated in figure 3b), adds relevant information and context to the decision process, and keeps track of our research question aimed at (balancing) priorities of local stakeholders. By complementing these quantitative values with possible qualitative preconceptions, the participants saw the impacts of their choices directly and could identify an energy system design best representing their preferences. After submitting their final energy system design, participants were also asked to choose up to three metrics (some chose more than three) that they prioritised for the energy system design — i.e. their stated priorities —, how much additional cost they were willing to accept to improve the representation of their priorities and provide written feedback to the process. Complete details on the interface, as well as the rationale for choosing the particular dimensions and metrics are described in Methods.
@@ 3 Cost-optimality paradigms cannot meet diverse expectations and technological perspectives
The interface outputs showcase the variety of proposed system designs and give insights about the acceptance of technologies, the importance of specific metrics as well as a partial rebuttal of cost-optimal designs. Through our exemplary work comprising 5% of the population (126 participants, demographics in Methods), we can also infer specific trends and patterns for system planning in Longyearbyen.
Figure 4: Combined strip- and boxplot of submitted results across the five dimensions. Stars represent the cost-optimal investment combination and one million NOK is approximately equivalent to 86,000 EUR or 90,500 USD in 2025. The min-max normalised median absolute deviations (MAD) quantify the distance between the median and the cost-optimum for the five dimensions, and are 0.06, 0.34, 0.11, 0.17 and 0.22, respectively. The central lines represent median values, the box limits represent the upper and lower quartiles, the whiskers extend from the box to the farthest data point lying within 1.5x the IQR. The outliers (along with all other data points) are shown in the underlying strip plot.
None of the technologies have consistent results with interquartile ranges (IQRs) of 10.9–36.8 percent of the decision interval. This is somewhat expected, as the design of the interface as well as the high slack level (125% more expensive systems are “near-optimal”) provide the participants with a high degree of freedom to select their preferred combination of technologies. In addition to the pre-computed 56,050 discrete energy system designs, the participants are also able to pick points anywhere within the feasible space, as the post-processing and interpolation within the interface (see Methods) increases the diversity of options.
@@
The resulting energy system designs of the participants show very high slack values (mean 81% and median 91%), raising questions about how well the additional cost is understood by the participants and who would bear that cost. This is striking given that almost half of the participants prioritised electricity prices when designing their system (more below). Such deviations from the cost optimum, however, are not unprecedented: indeed Fossas Tenas et al. [fossastenasEmpiricalEvidenceSlack2025] found median slack levels of above 40% in long-term transitions of European countries, with certain countries even exceeding the revealed slack levels of the participants. It needs to be noted that when the participants were explicitly asked about acceptable additional cost the results were much lower (mean and median values of 26% and 15% respectively) in the post-study questionnaire. This stark divide between revealed and stated preferences hints at the challenge to make results more tangible, but also shows that participants expect high performance on many priorities including affordability simultaneously.